# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of quantum complex flat connections at level-$\bm{k}$

### v1

@@ Abstract
A family of infinite-dimensional irreducible $\star$-representations on $\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{k}$ is defined for a quantum-deformed Lorentz algebra $U_{\mathbf{q}}(sl_{2})\otimes U_{\tilde{\mathbf{q}}}(sl_{2})$, where $\mathbf{q}=\exp[\frac{2\pi i}{k}(1+b^{2})]$ and $\tilde{\mathbf{q}}=\exp[\frac{2\pi i}{k}(1+b^{-2})]$ with $k\in\mathbb{Z}_{+}$ and $|b|=1$. The representations are constructed with the irreducible representation of quantum torus algebra at level-$k$, which is developed from the quantization of $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory. We study the Clebsch-Gordan decomposition of the tensor product representation, and we show that it reduces to the same problem as diagonalizing the complex Fenchel-Nielson length operators in quantizing $\mathrm{SL}(2,\mathbb{C})$ flat connections on 4-holed sphere. Finally, the spectral decomposition of the complex Fenchel-Nielson length operators results in the direct-integral representation of the Hilbert space $\mathcal{H}$, which we call the Fenchel-Nielson representation.
@@ 1 Introduction
The quantum Teichmüeller theory closely relates to the $\mathrm{SL}(2,\mathbb{R})$ Chern-Simons theory [8, 9]. There has been recent generalization in [10, 11, 12] to the Teichmüeller TQFT of integer level, which relates to the quantization of Chern-Simons theory with complex gauge group $\mathrm{SL}(2,\mathbb{C})$. The quantum $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory has the complex coupling constant $\hbar=\frac{2\pi i}{k}(1+b^{2}),\ \tilde{\hbar}=\frac{2\pi i}{k}(1+b^{-2})$ with $k\in\mathbb{Z},\ |b|=1$ (for one of the unitary branch [13]). The integer $k$ is called the level of Chern-Simons theory and relates to the integer level of the Teichmüeller TQFT. The quantization of $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory results in the Weyl algebra and quantum torus algebra at level-$k$, motivated by quantizing the Chern-Simon symplectic structure [12, 10]. The Weyl algebra has $\mathbf{q}=\exp(\hbar)$ and $\tilde{\mathbf{q}}=\exp(\tilde{\hbar})$, and the level-$k$ quantum torus algebra has $q=\exp(\hbar/2)$ and $\tilde{q}=\exp(\tilde{\hbar}/2)$. The Hilbert space carrying the representation is $\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{k}$, and the representation reduces to the representation in quantum Teichmüeller theory [5, 14] when $k=1$. The representation of quantum torus algebra at level-$k$ is reviewed in Section 2.
@@
Based on the representation of the quantum torus algebra, we construct a family of irreducible representation of $U_{\mathbf{q}}(sl_{2})\otimes U_{\tilde{\mathbf{q}}}(sl_{2})$ on $\mathcal{H}$. The star-structure on $U_{\mathbf{q}}(sl_{2})\otimes U_{\tilde{\mathbf{q}}}(sl_{2})$ is represented by the hermitian conjugate on $\mathcal{H}$. As the tensor product of two Hopf algebras, $U_{\mathbf{q}}(sl_{2})\otimes U_{\tilde{\mathbf{q}}}(sl_{2})$ has a well-defined Hopf algebra structure and can be understood as a quantum deformation of the Lorentz algebra. In particular, it is a generalization of the quantum Lorentz group with real $\mathbf{q}$ that is well-studied in the literature (see e.g.[15, 16]). The irreducible representations are parametrized by a continuous parameter $\mu_{L}\in\mathbb{R}$ and a discrete parameter $m_{L}\in\mathbb{Z}/k\mathbb{Z}$. They may be viewed as analog with the principle-series unitary representation of $\mathrm{SL}(2,\mathbb{C})$, although they are periodic in $m_{L}$.
@@
We study the Clebsch-Gordan decomposition of the tensor product representation (of $U_{\mathbf{q}}(sl_{2})\otimes U_{\tilde{\mathbf{q}}}(sl_{2})$) on $\mathcal{H}\otimes\mathcal{H}$, and we show that the result is a direct-integral of irreducible representations (see Section 3.2). The direct-integral is given by the spectral decomposition of the Casimir operator $Q_{21}$ of the tensor product representation, and we find the unitary transformation $\mathcal{U}_{21}$ as the Clebsch-Gordan map representing the co-multiplication.
@@
Interestingly $Q_{21}$ relates to the quantization of complex FN lengths for $\mathrm{SL}(2,\mathbb{C})$ flat connections on 4-holed sphere (see Section 4). As resulting from quantizing the $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory, the quantization of $\mathrm{SL}(2,\mathbb{C})$ flat connections on 4-holed sphere can be constructed based on the level-$k$ representation of quantum torus algebra. We focus on the case that $k=2N$ is even ($N$ is odd). In this case, we have the decomposition $\mathcal{H}\simeq\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ where each of ±$\mathcal{H}_{\pm}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$ carries the irreducible representation of the quantum algebra from quantizing the Fock-Goncharov (FG) coordinates of flat connections. The complex FN length that relates to $Q_{21}$ is given by the trace of holonomies around two holes. It turns out that the quantization of the complex FN length leads to the normal operators $\bm{L},\tilde{\bm{L}}=\bm{L}^{\dagger}$, and $\mathrm{id}_{\mathcal{H}}\otimes\bm{L}$ is represented the same as $Q_{21}$ on $\mathcal{H}\otimes\mathcal{H}$ up to a unitary transformation. Moreover, we show that the traces of S-cycle (enclosing the 1st and 2nd holes) and T-cycle (enclosing the 2nd and 3rd holes) holonomies are related by a unitary transformation, which is a realization of the A-move in the Moore-Seiberg groupoid (as a generalizaton from [6]).
@@
$\mathcal{H}\simeq\sum_{m_{r}\in\mathbb{Z}/k\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\rho(\mu_{r},m_{r})^{-1}\,\mathcal{H}_{\mu_{r},m_{r}},$ (1.1)
@@
where $\mathrm{d}\mu_{r}\rho(\mu_{r},m_{r})^{-1}$ is the spectral measure and each $\mathcal{H}_{\mu_{r},m_{r}}$ is 1-dimensional. The spectra of $\bm{L},\tilde{\bm{L}}$ are respectively $\ell(r)=r+r^{-1}$ and $\ell(r)^{*}$, where $r=\exp[\frac{2\pi i}{k}(-ib\mu_{r}-m_{r})]$. Due to the relation between $\bm{L}$ and $Q_{12}$, the direct-integral representation (1.1) also gives the Clebsch-Gordan decomposition for the tensor product representation of $U_{\mathbf{q}}(sl_{2})\otimes U_{\tilde{\mathbf{q}}}(sl_{2})$.
@@
The results from this work should have impact on the complex Chern-Simons theory at level-$k$ and its relation to quantum group and quantum Teichmüeller theory. For instance, although the quantum Lorentz group with real deformation and the representation theory has been widely studied, the generalization to complex $\mathbf{q}$ has not been studied in the literature before. We show in this paper that this generalization closely relates to the $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory with level-$k$  [^1] The quantum Lorentz group with real deformation relates to the $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory with $k=0$ [17].. As another aspect, given the relation between quantum Teichmüeller theory and Liouville conformal field theory, our study might point toward certain generalization of Liouville conformal field theory relating to the level-k, and this generalization might also relate to the boundary field theory of the $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory.
@@
In addition to the above, another motivation of this work is the potential application to the spinfoam model with cosmological constant in Loop Quantum Gravity [18, 19, 20]. The spinfoam model is formulated with $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory ($1/k$ is proportional to the absolute value of cosmological constant) with the special boundary condition called the simplicity constraint, which restricts the flat connections on 4-holed sphere to be SU(2) up to conjugation. Classically, the simplicity constraint is conveniently formulated in terms of the complex FN variables [21]. Therefore, it may be convenient to formulate the quantization of the simplicity constraint in the FN representation of quantum flat connections. The investigation on this perspective will be reported elsewhere.
@@
The structure of this paper is as follows: In Section 2, we review briefly the representation of quantum torus algebra at level-$k$ and set up some notations. In section 3, we construct the representations of $U_{\mathbf{q}}(sl_{2})\otimes U_{\tilde{\mathbf{q}}}(sl_{2})$ and discussion the Clebsch-Gordan decomposition of tensor product representation. In Section 4, we discuss the quantization of $\mathrm{SL}(2,\mathbb{C})$ flat connections on 4-holed sphere with a certain ideal triangulation, and we discuss the S-cycle and T-cycle trace operators and their unitary transformations. In Section 5, we compute the eigenvalue and distributional eigenstates of the trace operators and prove the direct-integral decomposition of the Hilbert space. In Section 6, we discuss the unitary transformation induced by changing ideal triangulation of the 4-holed sphere.
2 Quantum torus algebra and the representation at level-$k$
@@
The quantum torus algebra $\mathcal{O}_{q}(\mathbb{T}^{2})$ is spanned by Laurent polynomials of the symbols $\bm{x}_{\alpha,\beta}$ with $\alpha,\beta\in\mathbb{Z}$, satisfying the following relation
@@
where $\hbar\in\mathbb{C}$ is the quantum deformation parameter. We associated to $\mathcal{O}_{q}(\mathbb{T}^{2})$ the “anti-holomorphic counterpart” $\mathcal{O}_{\tilde{q}}(\mathbb{T}^{2})$ generated by $\tilde{\bm{x}}_{\alpha,\beta}$ with $\alpha,\beta\in\mathbb{Z}$, satisfying
@@
and $\mathcal{O}_{\tilde{q}}(\mathbb{T}^{2})$ commutes with $\mathcal{O}_{q}(\mathbb{T}^{2})$. The entire algebra is denoted by $\mathcal{A}_{\hbar}(\mathbb{T}^{2})=\mathcal{O}_{q}(\mathbb{T}^{2})\otimes\mathcal{O}_{\tilde{q}}(\mathbb{T}^{2})$. We can endow the algebra a $\star$-structure by
@@
It has been proposed in [12, 11] an infinite-dimensional unitary irreducible representation of $\mathcal{A}_{\hbar}(\mathbb{T}^{2})$ as the quantization of complex Chern-Simons theory. The Hilbert space carrying the representation is $\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{k}$, where a state is denoted by $f(\mu,m),$ $\mu\in\mathbb{R},$ $m\in\mathbb{Z}/k\mathbb{Z}$. The following operators are defined on $\mathcal{H}$
$\bm{\mu}f(\mu,m)$ $=\mu f(\mu,m),\qquad\bm{\nu}f(\mu,m)=-\frac{k}{2\pi i}\frac{\partial}{\partial\mu}f(\mu,m)$
$e^{\frac{2\pi i}{k}\bm{m}}f(\mu,m)$ $=e^{\frac{2\pi i}{k}m}f(\mu,m),\qquad e^{\frac{2\pi i}{k}\bm{n}}f(\mu,m)=f(\mu,m+1).$
@@
$[\bm{\mu},\bm{\nu}]=\frac{k}{2\pi i},\qquad e^{\frac{2\pi i}{k}\bm{n}}e^{\frac{2\pi i}{k}\bm{m}}=e^{\frac{2\pi i}{k}}e^{\frac{2\pi i}{k}\bm{m}}e^{\frac{2\pi i}{k}\bm{n}},\qquad[\bm{\nu},e^{\frac{2\pi i}{k}\bm{m}}]=[\bm{\mu},e^{\frac{2\pi i}{k}\bm{n}}]=0$ (2.3)
@@
$\bm{y}$ $=\exp\left[\frac{2\pi i}{k}(-ib\boldsymbol{\mu}-\bm{m})\right],$ $\qquad\tilde{\bm{y}}=\exp\left[\frac{2\pi i}{k}\left(-ib^{-1}\boldsymbol{\mu}+\bm{m}\right)\right],$ (2.4)
$\bm{x}$ $=\exp\left[\frac{2\pi i}{k}(-ib\boldsymbol{\nu}-\bm{n})\right],$ $\qquad\tilde{\bm{x}}=\exp\left[\frac{2\pi i}{k}\left(-ib^{-1}\boldsymbol{\nu}+\bm{n}\right)\right].$ (2.5)
@@
$\bm{y}f(\mu,m)$ $=e^{\frac{2\pi i}{k}(-ib{\mu}-{m})}f(\mu,m),\qquad$ $\bm{x}f(\mu,m)=f(\mu+ib,m-1),$ (2.6)
$\tilde{\bm{y}}f(\mu,m)$ $=e^{\frac{2\pi i}{k}(-ib^{-1}{\mu}+{m})}f(\mu,m),\qquad$ $\tilde{\bm{x}}f(\mu,m)=f(\mu+ib^{-1},m+1).$ (2.7)
@@
$\bm{x}\bm{y}=\mathbf{q}\bm{y}\bm{x},\qquad\tilde{\bm{x}}\tilde{\bm{y}}=\tilde{\mathbf{q}}\tilde{\bm{y}}\tilde{\bm{x}},\qquad\bm{x}\tilde{\bm{y}}=\tilde{\bm{y}}\bm{x},\qquad\tilde{\bm{x}}{\bm{y}}={\bm{y}}\tilde{\bm{x}}.$
@@
$\bm{Y}$ $=\frac{2\pi i}{k}(-ib\boldsymbol{\mu}-\bm{m}),\qquad\tilde{\bm{Y}}=\frac{2\pi i}{k}\left(-ib^{-1}\boldsymbol{\mu}+\bm{m}\right),$
$\bm{X}$ $=\frac{2\pi i}{k}(-ib\boldsymbol{\nu}-\bm{n}),\qquad\tilde{\bm{X}}=\frac{2\pi i}{k}\left(-ib^{-1}\boldsymbol{\nu}+\bm{n}\right),$
@@
The operators $\bm{x},\tilde{\bm{x}},\bm{y},\tilde{\bm{y}}$ are unbounded operators. The common domain $\mathfrak{D}$ of their Laurent polynomials contains $f(\mu,m)$ being entire functions in $\mu$ and satisfying
$e^{\alpha\frac{2\pi}{k}b\mu}f(\mu,m)\in L^{2}(\mathbb{R}),\qquad f(\mu+ib\alpha,m)\in L^{2}(\mathbb{R}),\qquad\forall\,m\in\mathbb{Z}/k\mathbb{Z},\quad\alpha\in\mathbb{Z}.$
The Hermite functions $e^{-\mu^{2}/2}H_{n}(\mu)$ , $n=1,\cdots,\infty$ satisfy all the requirements and span a dense domain in $L^{2}(\mathbb{R})$, so $\mathfrak{D}$ is dense in $\mathcal{H}$.
We denote by $\mathcal{L}(\mathfrak{D})$ the space of linear operators on $\mathfrak{D}$. The representation $\rho$: $\mathcal{A}_{\hbar}(\mathbb{T}^{2})\to\mathcal{L}(\mathfrak{D})$ is given by
$\rho:$ $\bm{x}_{\alpha,\beta}\mapsto e^{\alpha\bm{X}+\beta\bm{Y}}=q^{-\alpha\beta}\bm{x}^{\alpha}\bm{y}^{\beta},$ (2.8)
@@
The relations (2.1) and (2.2) are obtained by appyling the $(\mathbf{q},\tilde{\mathbf{q}})$-Weyl algebra. In the following, we often denote $\rho(\bm{x}_{\alpha,\beta}$) by $\bm{x}_{\alpha,\beta}$ for simplifying notations. The $\star$-stucture is represened by the Hermitian conjugate

### v2

# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of complex Chern-Simons theory at level-$\bm{N}$

@@ Abstract
A family of infinite-dimensional irreducible *-representations on $\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$ is defined for a quantum-deformed Lorentz algebra $\mathscr{U}_{\bf q}(sl_{2})\otimes\mathscr{U}_{\widetilde{\bf{q}}}(sl_{2})$, where $\mathbf{q}=\exp[\frac{\pi i}{N}(1+b^{2})]$ and $\tilde{\mathbf{q}}=\exp[\frac{\pi i}{N}(1+b^{-2})]$ with $N\in\mathbb{Z}_{+}$ and $|b|=1$. The representations are constructed with the irreducible representation of quantum torus algebra at level-$N$, which is developed from the quantization of $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory. We study the Clebsch-Gordan decomposition of the tensor product representation, and we show that it reduces to the same problem as diagonalizing the complex Fenchel-Nielson length operators in quantizing $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory on 4-holed sphere. Finally, we explicitly compute the spectral decomposition of the complex Fenchel-Nielson length operators and the corresponding direct-integral representation of the Hilbert space $\mathcal{H}$, which we call the Fenchel-Nielson representation.
@@ 1 Introduction
The quantum Teichmüeller theory closely relates to the $\mathrm{SL}(2,\mathbb{R})$ Chern-Simons theory [8, 9]. There has been recent generalization in [10, 11, 12] to the Teichmüeller TQFT of integer level, which relates to the quantization of Chern-Simons theory with complex gauge group $\mathrm{SL}(2,\mathbb{C})$. The quantum $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory has the complex coupling constant $N+is$ and $N-is$ where $N\in\mathbb{Z}_{+}$ is called the level of Chern-Simons theory, and $s\in\mathbb{R}$ corresponds to one of the unitary branch [13]. We use the parametrization
$is=N\frac{1-b^{2}}{1+b^{2}},\qquad|b|=1,\quad\mathrm{Re}(b)>0.$ (1.1)
and define $\hbar=\frac{\pi i}{N}(1+b^{2}),\ \tilde{\hbar}=\frac{\pi i}{N}(1+b^{-2})$. $N$ relates to the integer level of the Teichmüeller TQFT [11]. The quantization of $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory results in the Weyl algebra and quantum torus algebra at level-$N$, motivated by quantizing the Chern-Simon symplectic structure [12, 10]. The level-$N$ quantum torus algebra has $\mathbf{q}=\exp(\hbar)$ and $\tilde{\mathbf{q}}=\exp(\tilde{\hbar})$, and the Weyl algebra has $\mathbf{q}^{2}$ and $\tilde{\mathbf{q}}^{2}$. The Hilbert space carrying their representations is $\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$, and the representation reduces to the representation in quantum Teichmüeller theory [5, 14] when $N=1$. The representation of quantum torus algebra is reviewed in Section 2.
@@
Based on the representation of the quantum torus algebra, we construct a family of irreducible *-representation of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ on $\mathcal{H}$. The *-structure on $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ is represented by the hermitian conjugate on $\mathcal{H}$. As the tensor product of two Hopf algebras, $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ has a well-defined Hopf algebra structure and can be understood as a quantum deformation of the Lorentz algebra with complex deformation parameter $\mathbf{q},\tilde{\mathbf{q}}$. In particular, it is different from the quantum Lorentz group with real $q$ that is well-studied in the literature (see e.g. [15, 16]). The irreducible representations of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ constructed in this paper are parametrized by a continuous parameter $\mu_{L}\in\mathbb{R}$ and a discrete parameter $m_{L}\in\mathbb{Z}/k\mathbb{Z}$. They may be viewed as analog with the principle-series unitary representation of $\mathrm{SL}(2,\mathbb{C})$.
@@
We study the Clebsch-Gordan decomposition of the tensor product *-representation (of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$) on $\mathcal{H}\otimes\mathcal{H}$, and we show that the result is a direct-integral of irreducible *-representations (see Section 3.2). The direct-integral is given by the spectral decomposition of the Casimir operator $Q_{21}$ of the tensor product representation, and we find the unitary transformation $\mathcal{U}_{21}$ as the Clebsch-Gordan map representing the co-multiplication.
@@
Interestingly $Q_{21}$ closely relates to the quantization of complex FN lengths for $\mathrm{SL}(2,\mathbb{C})$ flat connections on 4-holed sphere (see Section 4). As resulting from quantizing the $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory, the quantization of $\mathrm{SL}(2,\mathbb{C})$ flat connections on 4-holed sphere can be constructed based on the level-$N$ representation of quantum torus algebra. The Hilbert space $\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$ carries the irreducible representation of the quantum algebra from quantizing the Fock-Goncharov (FG) coordinates of flat connections. The complex FN length that relates to $Q_{21}$ is given by the trace of holonomies around two holes. It turns out that the quantization of the complex FN length leads to the normal operators $\bm{L},\tilde{\bm{L}}=\bm{L}^{\dagger}$, and $\mathrm{id}_{\mathcal{H}}\otimes\bm{L}$ is unitary equivalent to $Q_{21}$ on $\mathcal{H}\otimes\mathcal{H}$. Moreover, we show that the traces of S-cycle (enclosing the 1st and 2nd holes) and T-cycle (enclosing the 2nd and 3rd holes) holonomies are related by a unitary transformation, which is a realization of the A-move in the Moore-Seiberg groupoid (as a generalizaton from [6]). In addition, we show that the quantization of holonomies’ traces on $\mathcal{H}$ gives a representation of the operator algebra from the skein quantization [17] of the flat connections.
@@
$\mathcal{H}\simeq\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\varrho(\mu_{r},m_{r})^{-1}\,\mathcal{H}_{\mu_{r},m_{r}},$ (1.2)
@@
where $\mathrm{d}\mu_{r}\varrho(\mu_{r},m_{r})^{-1}$ is the spectral measure and each $\mathcal{H}_{\mu_{r},m_{r}}$ is 1-dimensional. The spectra of $\bm{L},\tilde{\bm{L}}$ are respectively $\ell(r)=r+r^{-1}$ and $\ell(r)^{*}$, where $r=\exp[\frac{2\pi i}{N}(-ib\mu_{r}-m_{r})]$. Due to the relation between $\bm{L}$ and $Q_{12}$, the direct-integral representation (1.2) also gives the Clebsch-Gordan decomposition for the tensor product representation of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$.
@@
The results from this work should have impact on the complex Chern-Simons theory at level-$N$ and its relation to quantum group and quantum Teichmüeller theory. For instance, although the quantum Lorentz group with real deformation and the representation theory has been widely studied, the generalization to complex $\mathbf{q}$ has not been studied in the literature before. We show in this paper that this generalization closely relates to the $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory with level-$N$  [^1] The quantum Lorentz group with real deformation relates to the $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory with $N=0$ [18].. As another aspect, given the relation between quantum Teichmüeller theory and Liouville conformal field theory, our study might point toward certain generalization of Liouville conformal field theory relating to the level-k, and this generalization might also relate to the boundary field theory of the $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory.
@@
In addition to the above, another motivation of this work is the potential application to the spinfoam model with cosmological constant [19, 20, 21]. The spinfoam model is formulated with $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory ($1/N$ is proportional to the absolute value of cosmological constant) with the special boundary condition called the simplicity constraint, which restricts the flat connections on 4-holed sphere to be SU(2) up to conjugation. Classically, the simplicity constraint is conveniently formulated in terms of the complex FN variables [22]. Therefore, it may be convenient to formulate the quantization of the simplicity constraint in the FN representation of quantum flat connections. The investigation on this perspective will be reported elsewhere.
@@
The structure of this paper is as follows: In Section 2, we review briefly the representation of quantum torus algebra at level-$k$ and set up some notations. In section 3, we construct the representations of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ and discussion the Clebsch-Gordan decomposition of tensor product representation. In Section 4, we discuss the quantization of $\mathrm{SL}(2,\mathbb{C})$ flat connections on 4-holed sphere with a certain ideal triangulation, the relation with the skein quantization, the S-cycle and T-cycle trace operators and their unitary transformations. In Section 5, we compute the eigenvalue and distributional eigenstates of the trace operators and prove the direct-integral decomposition of the Hilbert space. In Section 6, we discuss the unitary transformation induced by changing ideal triangulation of the 4-holed sphere.
2 Quantum torus algebra and the representation at level $k=2N$
@@
The quantum torus algebra $\mathcal{O}_{q}$ is spanned by Laurent polynomials of the symbols $\bm{x}_{\alpha,\beta}$ with $\alpha,\beta\in\mathbb{Z}$, satisfying the following relation
@@
where $\hbar\in\mathbb{C}$ is the quantum deformation parameter. We associated to $\mathcal{O}_{q}$ the “anti-holomorphic counterpart” $\mathcal{O}_{\tilde{q}}$ generated by $\tilde{\bm{x}}_{\alpha,\beta}$ with $\alpha,\beta\in\mathbb{Z}$, satisfying
@@
and $\mathcal{O}_{\tilde{q}}$ commutes with $\mathcal{O}_{q}$. The entire algebra is denoted by $\mathcal{A}_{\hbar}=\mathcal{O}_{q}\otimes\mathcal{O}_{\tilde{q}}$. We can endow the algebra a $\star$-structure by
@@
It has been proposed in [12, 11] an infinite-dimensional unitary irreducible representation of $\mathcal{A}_{\hbar}$ as the quantization of complex Chern-Simons theory. The Hilbert space carrying the representation is $\mathcal{H}_{0}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{k}$, where a state is denoted by $f(\mu_{0},m_{0}),$ $\mu_{0}\in\mathbb{R},$ $m_{0}\in\mathbb{Z}/k\mathbb{Z}$. The following operators are defined on $\mathcal{H}_{0}$
$\bm{\mu}_{0}f(\mu_{0},m_{0})$ $=\mu_{0}f(\mu_{0},m_{0}),\qquad\bm{\nu}_{0}f(\mu_{0},m_{0})=-\frac{k}{2\pi i}\frac{\partial}{\partial\mu_{0}}f(\mu_{0},m_{0})$
$e^{\frac{2\pi i}{k}\bm{m}_{0}}f(\mu_{0},m_{0})$ $=e^{\frac{2\pi i}{k}m_{0}}f(\mu_{0},m_{0}),\qquad e^{\frac{2\pi i}{k}\bm{n_{0}}}f(\mu_{0},m_{0})=f(\mu_{0},m_{0}+1).$
@@
$[\bm{\mu}_{0},\bm{\nu}_{0}]=\frac{k}{2\pi i},\qquad e^{\frac{2\pi i}{k}\bm{n}_{0}}e^{\frac{2\pi i}{k}\bm{m}_{0}}=e^{\frac{2\pi i}{k}}e^{\frac{2\pi i}{k}\bm{m}_{0}}e^{\frac{2\pi i}{k}\bm{n}_{0}},\qquad[\bm{\nu}_{0},e^{\frac{2\pi i}{k}\bm{m}_{0}}]=[\bm{\mu}_{0},e^{\frac{2\pi i}{k}\bm{n}_{0}}]=0$ (2.3)
@@
$\bm{y}$ $=$ $\exp\left[\frac{2\pi i}{k}(-ib\boldsymbol{\mu}_{0}-\bm{m}_{0})\right],\qquad\tilde{\bm{y}}=\exp\left[\frac{2\pi i}{k}\left(-ib^{-1}\boldsymbol{\mu}_{0}+\bm{m}_{0}\right)\right],$ (2.4)
$\bm{x}$ $=$ $\exp\left[\frac{2\pi i}{k}(-ib\boldsymbol{\nu}_{0}-\bm{n}_{0})\right],\qquad\tilde{\bm{x}}=\exp\left[\frac{2\pi i}{k}\left(-ib^{-1}\boldsymbol{\nu}_{0}+\bm{n}_{0}\right)\right].$ (2.5)
@@
$\bm{y}f(\mu_{0},m_{0})$ $=$ $e^{\frac{2\pi i}{k}(-ib{\mu}_{0}-{m}_{0})}f(\mu_{0},m_{0}),\qquad\bm{x}f(\mu_{0},m_{0})=f(\mu_{0}+ib,m_{0}-1),$ (2.6)
$\tilde{\bm{y}}f(\mu_{0},m_{0})$ $=$ $e^{\frac{2\pi i}{k}(-ib^{-1}{\mu}_{0}+{m}_{0})}f(\mu_{0},m_{0}),\qquad\tilde{\bm{x}}f(\mu_{0},m_{0})=f(\mu_{0}+ib^{-1},m_{0}+1).$ (2.7)
@@
$\bm{x}\bm{y}=\mathbf{q}\bm{y}\bm{x},\qquad\tilde{\bm{x}}\tilde{\bm{y}}=\tilde{\mathbf{q}}\tilde{\bm{y}}\tilde{\bm{x}},\qquad\bm{x}\tilde{\bm{y}}=\tilde{\bm{y}}\bm{x},\qquad\tilde{\bm{x}}{\bm{y}}={\bm{y}}\tilde{\bm{x}}.$ (2.8)
@@
$\bm{Y}$ $=\frac{2\pi i}{k}(-ib\boldsymbol{\mu}_{0}-\bm{m}_{0}),\qquad\tilde{\bm{Y}}=\frac{2\pi i}{k}\left(-ib^{-1}\boldsymbol{\mu}_{0}+\bm{m}_{0}\right),$
$\bm{X}$ $=\frac{2\pi i}{k}(-ib\boldsymbol{\nu}_{0}-\bm{n}_{0}),\qquad\tilde{\bm{X}}=\frac{2\pi i}{k}\left(-ib^{-1}\boldsymbol{\nu}_{0}+\bm{n}_{0}\right),$
@@
The operators $\bm{x},\tilde{\bm{x}},\bm{y},\tilde{\bm{y}}$ are unbounded operators. The common domain $\mathfrak{D}_{0}$ of their Laurent polynomials contains $f(\mu_{0},m_{0})$ being entire functions in $\mu_{0}$ and satisfying
$e^{\alpha_{1}\frac{2\pi}{k}b\mu_{0}}e^{\alpha_{2}\frac{2\pi}{k}b^{-1}\mu_{0}}f(\mu_{0}+ib\beta_{1}+ib^{-1}\beta_{2},m_{0})\in L^{2}(\mathbb{R}),\qquad\forall\,m_{0}\in\mathbb{Z}/k\mathbb{Z},\quad\alpha_{i},\beta_{i}\in\mathbb{Z}.$ (2.9)
The Hermite functions $e^{-\mu_{0}^{2}/2}H_{n}(\mu_{0})$ , $n=1,\cdots,\infty$ satisfy all the requirements and span a dense domain in $L^{2}(\mathbb{R})$, so $\mathfrak{D}_{0}$ is dense in $\mathcal{H}_{0}$.
We denote by $\mathcal{L}(\mathfrak{D}_{0})$ the space of linear operators on $\mathfrak{D}_{0}$. The representation $\rho_{0}$: $\mathcal{A}_{\hbar}\to\mathcal{L}(\mathfrak{D}_{0})$ is given by
$\rho_{0}:$ $\bm{x}_{\alpha,\beta}\mapsto e^{\alpha\bm{X}+\beta\bm{Y}}=q^{-\alpha\beta}\bm{x}^{\alpha}\bm{y}^{\beta},$ (2.10)
@@
The relations (2.1) and (2.2) are obtained by applying the $(\mathbf{q},\tilde{\mathbf{q}})$-Weyl algebra. In the following, we often denote $\rho(\bm{x}_{\alpha,\beta}$) by $\bm{x}_{\alpha,\beta}$ for simplifying notations. The $\star$-stucture is represented by the Hermitian conjugate